OpenStudy (anonymous):

graph the function and state the vertex, axis of symmetry, and the intercepts, if any. f(x)=x^2-6x+8

OpenStudy (anonymous):

This is easy.

OpenStudy (anonymous):

First you have to find the "zeroes" or the x-intercepts.

OpenStudy (anonymous):

so do i replace all x's with 0?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You have to replace Y AS 0, because the x-intercept is what x is when y = 0.

OpenStudy (anonymous):

So we have 0 = x^2 - 6x + 8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now we have to factor.

OpenStudy (anonymous):

do you have to get rid of the ^2?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

im sorry math is not my strong suit, this stuff really confuses me =(

OpenStudy (anonymous):

Well, later on we do.

OpenStudy (anonymous):

It's alright.

OpenStudy (anonymous):

So we have to find out the squares of the first and last terms, do you know what they are?

OpenStudy (anonymous):

Well we get x from the first term, and we use 8 as our last term.

OpenStudy (anonymous):

So 8 * x = 8x so we need to find out what the factors of 8 are that add up to the MIDDLE TERM 6x.

OpenStudy (anonymous):

Can you state some factors of 8?

OpenStudy (anonymous):

1,2,4,8

OpenStudy (anonymous):

Right, so let's try 1 and 8.

OpenStudy (anonymous):

Remember, they both have to add up to 6 and multiply to 8.

OpenStudy (anonymous):

1 x 8 = 8. But 1 + 8 = 9. We're looking for a 6, not a 9.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2 and 4

OpenStudy (anonymous):

Yes, you got it!

OpenStudy (anonymous):

So we replace 6x with 4x and 2x.

OpenStudy (anonymous):

0= x^2-(4x+2x)+8?

OpenStudy (anonymous):

Uh-oh

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I made a mistake, sorry. We need 2 things that add up to -6 and multiply to 8.

OpenStudy (anonymous):

so -2 and 4

OpenStudy (anonymous):

-2 and -4*

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So we have 0= x^2 - 4x - 2x + 8?

OpenStudy (anonymous):

Now we're going to separate the first two terms and the last two terms with a parenthesis:

OpenStudy (anonymous):

0 - (x^2 - 4x) - (2x + 8)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We have to find out the GCF or greatest common factor of each parenthesis, can you tell what the GCF of the first parenthesis is?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

No, look again. Remember it can be a variable.

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So we're going to work backwards with the first parenthesis: x(x^2 - 4x) We are DIVIDING x by each term, since we're working backwards

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So basically we're removing 1 x from all the parenthesis: x(x^2 - 4x) Removing 1 x from x^2 would give us a regular x. x(x - 4x) And removing x from -4x will give us a -4. x(x - 4) So we finished our first parenthesis.

OpenStudy (anonymous):

Can you tell me the GCF of the 2nd parenthesis?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so it would be 2(2x+4)?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

2(x+4)*

OpenStudy (anonymous):

You made a mistake, remember that your 2x is a NEGATIVE 2x.

OpenStudy (anonymous):

so 2(-x+4)?

OpenStudy (anonymous):

No, your GCF will be a -2.

OpenStudy (anonymous):

You want to keep x positive.

OpenStudy (anonymous):

-2(x+4)

OpenStudy (anonymous):

Yes, almost perfect. Remember 8 / -2 = -4, not 4.

OpenStudy (anonymous):

-2(x-4)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So we have 0 = x(x - 4) - 2(x - 4) We can combine x - 4, because they're the same(If these terms aren't the same, then you know you made a mistake). And we have to combine the terms outside of both parentheses. So we have (x - 2)(x - 4) Do you understand how I got (x - 2)?

OpenStudy (anonymous):

no =(

OpenStudy (anonymous):

oooh wait! the x and -2 outside of the parentheses!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You need to combine those.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So now that we figured out that form we can finish the problem. 0 = (x - 2)(x - 4)

OpenStudy (anonymous):

We have to break each parenthesis down and add a '= 0' to it. So for the first parenthesis: 0 = x - 2 Now we just figure out what x is by adding 2 to both sides. 2 = x x = 2 Can you figure out the 2nd parenthesis out?

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

Yes, so we have: x = 2 and x = 4. We know that the parabola crosses the x-axis at 2 and 4.

OpenStudy (anonymous):

Those are your x-intercepts.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The y-intercept is what y is when x = 0. So we can go back to the original equation and change all the x's to 0. f(x) = (0)^2 - 6(0) + 8 f(x) = 0 - 0 + 8 f(x) = 0 + 8 f(x) = 8 So our y-intercept is 8.

OpenStudy (anonymous):

would it be (0,8) for the y intercept and (2,0) (4,0) for x?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok =)

OpenStudy (anonymous):

Now we have to find the vertex.

OpenStudy (anonymous):

The formula for the vertex is -b/2a.

OpenStudy (anonymous):

Your equation is in the form of a^2 - bx + c So what is 'a' and 'b'?

OpenStudy (anonymous):

x and -6

OpenStudy (anonymous):

Not quite, the "a" has to be a number.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

x by itself can also be what?

OpenStudy (anonymous):

Yes, that's it.

OpenStudy (anonymous):

So we have 1 and -6. We can plug these into our equation: -1/2(-6) Can you figure it out?

OpenStudy (anonymous):

My bad it's: -6/2(-1)

OpenStudy (anonymous):

6/-2

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Ugh, I made another mistake it's 6/2(1) 6/2 3

OpenStudy (anonymous):

3 is the x-value of our vertex. We can plug it into our equation to get the y-intercept.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

f(x) = x^2 - 6x + 8 f(x) = (3)^2 - 6(3) + 8

OpenStudy (anonymous):

f(x) = (3)^2 - 6(3) + 8 f(x) = 9 - 18 + 8 f(x) = -9 + 8 f(x) = -1 So the y-value of our vertex is -1.

OpenStudy (anonymous):

yes i got -1

OpenStudy (anonymous):

vertex is (3,-1) ?

OpenStudy (anonymous):

So our vertex is: (3, -1). Our y-intercept is (0, 8), and our x-intercepts are (2, 0) and (4, 0).

OpenStudy (anonymous):

Yes, it's (3, -1)

OpenStudy (anonymous):

The axis of symmetry will be our x-value, or 3.

OpenStudy (anonymous):

oh ok awesome!

OpenStudy (anonymous):

That's everything, that's how you figure it out by looking at it from our equation. You can graph it here: https://www.desmos.com/calculator And you will get all the instant answers.

OpenStudy (anonymous):

thank you so much!!! =)